OpenStudy (anonymous):

Two positive integers have a product of 176. One integer is 5 less than the other integer. Which equation can be used to find the value of x, the greater integer? x2 + 5 = 176 x(x + 5) = 176 x(x – 5) = 176 x2 – 5 = 176

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

are you missing an x in the original?

OpenStudy (anonymous):

What Do u Mean

OpenStudy (triciaal):

by the way, please read the post to the last question already closed the answer choice was written wrong

OpenStudy (anonymous):

ok and how do i do this

OpenStudy (triciaal):

x(x – 5) = 176 product so multiply x is larger x-5 is the other less mean you subtract

OpenStudy (triciaal):

ok?

OpenStudy (anonymous):

Yea So It Would Be D

OpenStudy (anonymous):

@triciaal ^

OpenStudy (anonymous):

@ganeshie8

OpenStudy (triciaal):

x(x – 5) = 176 this is option C

OpenStudy (triciaal):

the larger number X multiplied by the smaller (x-5) = 176

OpenStudy (anonymous):

iSee Now

OpenStudy (triciaal):

great